(10x15)+(6x+25)=180
Answer by ikdeep(189)

(10x15)+(6x+25)=180
first solve the brackets and we get
150 + (6x+25) = 180
subtract 150 from both sides and we get
6x+25 = 30
now subtarct 25 from both sides and we get ...
6x = 5
divide both sides by 6 adn we get ...
x = 5/6
to verify put x=5/6 adn we get LHS=RHS then our answer is correct ...
hope this will help u
feel free to revert